# Dependency-Aware Semi-Structured Sparsity of GLU Variants in Large Language Models

### v1

@@ Appendix A Model Configurations
Here are the configurations of the models used in the paper. We don’t use LLaMA2-34B since it was not released. ReluLLaMA uses the same configuration as LLaMA2, the only difference is the activation function. Here is the link to ReluLLaMA:
ReluLLaMA-7B: https://huggingface.co/SparseLLM/ReluLLaMA-7B
ReluLLaMA-13B: https://huggingface.co/SparseLLM/ReluLLaMA-13B
Table 7: Model configurations of Llama2 and Mistral models.
@@
Appendix B Activation Outlier in SwiGLU
Figure 4: Layerwise Activation Outlier distribution of LLaMA2-7B
In Section 2.3, we show the importance of augmenting input magnitude is useful. In this part, we target to identify the existence of outliers in the intermediate activations of SwiGLU. Directly analyzing it like model hidden states Dettmers et al. (2022); Xiao et al. (2023) is challenging since it demonstrates high variance. (Ashkboos et al., 2023). Wanda (Sun et al., 2023) shows $\ell_{2}$ norm can be a metric for evaluating activations, it shows effectiveness in Down-Proj pruning. Motivated by Yin et al. (2023), we define the outlier in MLP intermediate activations by identifying activations whose $\ell_{2}$ norm magnitude is $T$ times greater than the average MLP activation norm in the corresponding layer. We set $T$ as an empirical value 7. The layerwise distribution of the SwiGLU intermediate activation outlier is shown in Figure 4.
Appendix C GeLU-based MLP Input projection
In the older MLP module, unlike GLU variants, there are only two projection matrices. We investigate magnitude pruning in different granularity for the MLP input projection matrix. We use the Falcon-7B model, which uses GeLU-based MLP. ”Layer Magnitude” means we group weights in the entire weight matrix using magnitude pruning. In Table 8, we surprisingly find the input-balanced magnitude pruning achieves much worse results. This is in contrast to the results in GLU variants. The results show the huge significance of input activations for GeLU-based input projection pruning. Since our DaSS method needs to group weights per input, our method is less applicable to older MLP modules. However, GeLU-based MLP is rarely used in the current state-of-the-art LLMs.
Sparsity 50% 60% 70%
Wanda 7.17 8.72 18.95
Output-balanced Magnitude 8.46 15.27 498.77
Input-balanced Magnitude 1708 10587 13706
Layer Magnitude 21.82 1895 12896

### v2

@@ A.1 Model Configurations
Table 11 is the configurations of the models used in the paper. We did not use LLaMA2-34B as it was not released. ReluLLaMA uses the same configuration as LLaMA2, with the only difference being the activation function. The links to the ReluLLaMA models are provided below:
A.2 Robustness to calibration samples.
Figure 2: Robustness to calibration samples.
A.3 GeLU-based MLP Input projection
In the older MLP module, unlike GLU variants, there are only two projection matrices. We investigate magnitude pruning in different granularity for the MLP input projection matrix. We use the Falcon-7B model, which uses GeLU-based MLP. "Layer Magnitude" means we group weights in the entire weight matrix using magnitude pruning. In Table 7, we surprisingly find the input-balanced magnitude pruning achieves much worse results. This is in contrast to the results in GLU variants. The results show the huge significance of input activations for GeLU-based input projection pruning. Since our DaSS method needs to group weights per input, our method is less applicable to older MLP modules. However, GeLU-based MLP is rarely used in the current state-of-the-art LLMs.
Sparsity 50% 60% 70%
Wanda 7.17 8.72 18.95
Output-balanced Magnitude 8.46 15.27 498.77
Input-balanced Magnitude 1708 10587 13706
Layer Magnitude 21.82 1895 12896
A.4 Combination with Quantization
Method SaprseGPT Wanda DaSS
50% sparsity 6.38 6.50 6.44
w/ 4-bit AWQ 6.51 6.63 6.57
Table 8: The Wikitext perplexity results of LLaMA2-7B with 50% sparsity and 4 bit quantization
Pruning and quantization, traditionally seen as distinct model compression techniques, are not mutually exclusive and can be effectively integrated to enhance overall efficiency Han et al. (2015a); Frantar and Alistarh (2023); Agarwalla et al. (2024). Here we test the combination of DaSS with 4-bit AWQ Lin et al. (2024) for compressing LlaMA2-7B. As shown in Table 8, the performance of all the pruned models just drops slightly after 4-bit quantization. DaSS still outperforms wanda after 4-bit quantization.
A.5 MMLU Results
Concurrent work (Gromov et al., 2024) indicates the unusual behavior of LLMs in performing Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2021) tasks. We use Chain-of-Thought Hub Fu et al. (2023) which is based on the official implementation of MMLU Hendrycks et al. (2021). The MMLU encompasses 57 tasks, spanning from STEM, Humanities, Social Sciences, among others and we report the mean accuracy of 57 tasks. We test the performance of pruned LlaMA2-70B model in MMLU task.
Sparsity Methods MMLU (5 shot)
Dense - 69.10
4:8 SparseGPT 60.24
Wanda 59.69
DaSS 60.86
DaSS+skip 1/4 65.82
2:4 SparseGPT 56.99
Wanda 56.41
DaSS 57.53
DaSS+skip 1/4 64.36
50% SparseGPT 64.52
Wanda 62.72
DaSS 63.27
DaSS+skip 1/4 66.81
Table 9: MMLU scores for different sparsity methods and levels
From Table 9, we observe that the improvement of DaSS over Wanda becomes more pronounced in the challenging MMLU task, where it achieves an increase in accuracy of 1.17 and 1.12 for the 4:8 and 2:4 sparsity patterns, respectively. In accordance with the observations with Jaiswal et al. (2023), we see a considerable performance degradation in knowledge-intensive MMLU task, in which only unstructured 50% sparsity models outperform the dense LLaMA2-34B model. As GPU plays a more important role in larger model inference, it is essential to improve the performance of pruned models in hardware-friendly N:M sparsity. Uniform reduction of the overall sparsity level is not feasible for N:M sparsity, Frantar and Alistarh (2023) suggests a specific subset of layers can be chosen for full N:M sparsification. Here we skip pruning 1/4 consecutive layers (20 layers) and result in a final 37.5% sparsity ratio. We use the first 10 tasks of MMLU to study pruning sensitivity. We divide the model into 4 consecutive parts and study skipping pruning each part. As shown in Table 10, the earlier layers are more sensitive than the later ones in knowledge-intensitive tasks, which is contradictory to the findings in Frantar and Alistarh (2023) using perplexity, and align with the findings of Gromov et al. (2024).
Start Index 0 10 20 40 60 Dense
Acc (%) 60.12 61.72 59.75 56.54 56.83 64.86
@@ A.5 MMLU Results
Gemma-7B 7B 28 3072 24576 16 16
Table 11: Model configurations of Llama2, Gemma and Mistral models.
Concurrent work Gromov et al. (2024) propose an effective method to remove half of the layers in LlaMA2-70B with small accuracy loss for MMLU tasks, suggesting the unusual behavior of LLMs in performing MMLU tasks. Their work is focused more on task-specific pruning, while our work emphasizes preserving the general ability of LLMs to perform various tasks. Our weight pruning method is fundamentally orthogonal to layer pruning in principle.